OpenStudy (anonymous):

1. Write the slope intercept form of the equation of the line that passes through (4,2) with a slope of 3. 2. When graphing a point for (-1,4) you would - move 1 right and 4 down - move 1 right and 4 up - move 1 left and 4 up - move 1 left and 4 down

OpenStudy (anonymous):

use this equation f(x)-f(a)=k(x-a)

OpenStudy (anonymous):

k=3 a=4 and f(a)=2

OpenStudy (anonymous):

so you should get f(x)=3(x-4)+2

OpenStudy (anonymous):

you can distribute that 3 into the parenthesis or leave it as is, doesn't really matter

OpenStudy (anonymous):

the answer choices I have are: - y = 3x + 4 - y = 4x + 3 - y = 2 = 3 ( x - 4) - y = 3x - 10 @jrn1007

OpenStudy (anonymous):

that last one

OpenStudy (anonymous):

@jrn1007 thanks :) and can you help me with the other question too?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

when graphing you would move to the right 1 and up 4

OpenStudy (anonymous):

@jrn1007 which of the following is the equation of the vertical line that passes through (4,6)? - x = 4 - y = 6 - y = 4 = x = 6 When graphing a point for (-1,4), you would: - move 1 right and 4 down - move 1 right and 4 up - move 1 left and 4 up - move 1 left and 4 down

OpenStudy (anonymous):

it is x=4

OpenStudy (anonymous):

notice they said vertical line so the only x value you have is point on the axis where the vertical line passes through

OpenStudy (anonymous):

@jrn1007 thank you so much by the way you are saving my life! my teacher always hands out these really complicated assignment sheets and never explains :)

OpenStudy (anonymous):

so the answer is move 1 left and 4 up? @jrn1007

OpenStudy (anonymous):

sure fan me and i'll fan you, hit me up anytime you need help... I can help all the way up through multi var calc

OpenStudy (anonymous):

thanks :) ! and the answer to the last question is "move 1 left and 4 up"? @jrn1007

OpenStudy (anonymous):

yes left 1 and up 4

OpenStudy (anonymous):

@jrn1007 can i ask two last questions?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

1. What is the range of this set of ordered pairs? {(5,2), (-5,0),(6,2),(2,7)} - {0, 2, 7} - {-5,0,2,5,6,7} - {-5,2,5,6} @jrn1007

OpenStudy (anonymous):

best way to do this is create a simple x,y plane, and plot the 4 points. then look only at the y values of those points. hold one sec while i do this.

OpenStudy (anonymous):

its the first one, (0,2,7)

OpenStudy (anonymous):

@jrn1007 ok thanks so much for all your help!

OpenStudy (anonymous):

when you have an ordered pair such as (a,b) where a and b are numbers the a corresponds to x values and b corresponds to the y values. so domain entails only looking at x-values and range entails only looking at the y-values. You're welcome!

OpenStudy (anonymous):

@jrn1007 sorry but I was just doing my homework and I found another problem I don't really understand Which of the following is the equation for the horizontal line that passes through (-1, -3)? - x= -1 - x = -3 - y - -3 - y = -1

OpenStudy (anonymous):

ok cool, notice that it says horizontal line, so there should be only ONE value of y and a whole domain of x's. so we would need one y value and that's the one they gave us. so y=-3 is our horizontal line

OpenStudy (anonymous):

@jrn1007 thanks!

OpenStudy (anonymous):

no prob! oh and if you ever need an equation for a vertical line it should be only ONE value of x and a whole range of y's, sort of the opposite. so our vertical line would be something like x=whatever number